# Diffusion Models for Time Series Forecasting: A Survey

### v1

@@ 3 Conditions for TSF Diffusion Models
Figure 2: Overview of Diffusion Models for TSF. Each approach is categorized by its conditional source and condition integration process. To the best of our knowledge, there is no study that uses a diffusion-centric model to encode multimodal data for TSF.
TSF with diffusion models is a conditional generation task; therefore, the design of conditioning information is critical to the forecasting performance. Especially for diffusion-based TSF approaches that use historical data as conditions, how much useful information is extracted from the historical data and the richness of the historical data are critical for the sequence modeling. This characteristic motivates diffusion-based TSF approaches to improve and expand their conditioning source. Therefore, we categorize existing research into two primary groups based on their historical conditioning sources: historical time series conditions and multimodal conditions. For historical time series conditions, we analyze how existing approaches explicitly incorporate specific temporal priors into the condition. This integration guides the sampling process towards generating accurate future time series consistent with underlying temporal patterns. For multimodal conditioning sources, we detail current diffusion-based multimodal time series approaches. These approaches incorporate non-time-series modality information from distinct perspectives, each possessing unique characteristics. The following sections elaborate on these different categories of studies, explaining their methodologies and characteristics.
@@ 3.1 Conditioning on Historical Time Series
Generally, the raw historical time series itself directly serves as the condition for diffusion models without requiring additional preprocessing. Most studies in the field utilize this straightforward approach, primarily focusing on improving the architecture of the diffusion model. Beyond using the raw series, applying specific time series processing techniques and transformations enables the extraction and explicit representation of more temporal information, thereby incorporating domain-specific prior knowledge as conditioning signals. For instance, decomposition techniques [49] separate the time series into distinct components, such as trend, seasonality, and residuals. These components capture different aspects of the temporal dynamics, enhancing models’ interpretability and simplifying the learning task by isolating underlying patterns. Consequently, approaches such as mr-Diff [50] and FDF [51] decompose the historical time series into trend and seasonal residual terms before diffusion. This decomposition explicitly injects disentangled temporal information as hierarchical conditions into the reverse process. Furthermore, time series inherently possess multi-scale characteristics, where different temporal scales reveal varying degrees of trend evolution and cyclical behavior. MG-TSD [52] and mr-Diff [50] leverage this property by employing representations of the series at multiple scales as conditions during the reverse process. This alignment connects the intermediate diffusion states to different granularity levels, mirroring the coarse-to-fine refinement in time series generation. Beyond these time-domain techniques, frequency-domain transformations convert the time-evolving sequence values into sparse spectral representations. Diffusion-TS [53] and FALDA [54] adopt this strategy, conditioning the diffusion process on both the original historical series and its frequency spectrum. This dual-domain conditioning supplies complementary perspectives on the sequence modeling. Other novel approaches include retrieval augmentation. RATD [55] integrates a mechanism inspired by the natural language processing domain, retrieving the most similar historical sequence segments from the training set as reference time series. Combining these retrieved references with the original time series provides conditions that help the model learn highly repetitive intrinsic temporal patterns effectively. In another direction of these approaches, LDT [46] employs a variational autoencoder to transform the time series into a compressed latent representation before diffusion. This latent space transformation automatically extracts salient features for conditioning while simultaneously reducing computational costs associated with high-dimensional time series.
@@ 3.2 Conditioning on Multimodal Data
Beyond preprocessing and transformation techniques that extract useful information from historical time series as domain-specific priors, incorporating multimodal data as additional condition sources provides richer conditioning information. Multimodal approaches [56] gain increasing attention in the TSF domain. Meanwhile, diffusion-based TSF approaches also integrate additional modal information as complementary conditions. For instance, LDM4TS [57] applies complementary vision transformation to convert time series into multi-view images. It also constructs textual descriptions for the time series using predefined templates filled with the statistical features of the time series. These generated images and text descriptions serve as multimodal conditions injected into the reverse process. They guide the diffusion model to predict future time series values based on both visual patterns and semantic text features. Crucially, LDM4TS [57] utilizes only the original time series data itself to create these modalities. This self-contained nature allows straightforward application to existing datasets. Furthermore, the MCD-TSF [58] approach incorporates external news article text recorded concurrently with the target time series. This external text potentially contains information absent from the series data itself. Such external multimodal data offers valuable supplementary conditioning, particularly in forecasting scenarios with frequent external interventions. Conversely, approaches relying on external data impose mandatory external data requirements on time series datasets, which may restrict their applicability across different domains or data availability contexts.
@@ 4 Condition Integration
In addition to the source used for conditioning, diffusion-based TSF approaches also differ in how they integrate conditioning information into the generative diffusion process. Different from image-to-text or image-to-image generation that focus on transferring features and styles, time series forecasting instead extends historical data into future sequences, so the condition integration mechanism is crucial for ensuring coherent evolution from past to future. Therefore, we classify existing condition integration mechanisms into two primary categories, feature-centric and diffusion-centric approaches. Feature-centric strategies [35, 59, 60] employ the standard diffusion model. These approaches rely on a well-designed network to extract useful features from the historical conditioning data. These features then guide each denoising step in the reverse diffusion process, ensuring that the generated sequence remains temporally consistent with its historical context. Consequently, the condition acts primarily as input to the learned denoising function. Diffusion-centric approaches [61, 62, 63] propose modifications to the diffusion process. Such techniques incorporate historical data priors directly into both forward and reverse diffusion trajectories, enabling each step to adapt dynamically to past information. By conditioning adjustments on the preceding context, they deliver finer-grained control over the generative process. This dynamic mechanism adapts diffusion dynamics based on the specific temporal features extracted from the historical data. An overview of the diffusion-based TSF approaches and their categorization by conditioning source and integration process is illustrated in Figure 2. The subsequent sections explain these various study categories, detailing their methodologies and characteristics.
@@ 4.1 Feature-centric Condition
The feature-centric approaches [35, 59, 60] retain the standard diffusion model’s forward noising and reverse denoising processes. These approaches focus on designing an effective feature extractor that is capable of extracting time-dependent features or modeling time-series uncertainty from historical data, which is formulated as
$p_{\theta}(\mathbf{x}_{1:H}^{t-1}|\mathbf{x}_{1:H}^{t},\mathbf{x}_{-L+1:0})=\mathcal{N}(\mathbf{x}_{1:H}^{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{1:H}^{t},t,\phi(\mathbf{x}_{-L+1:0})),\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{1:H}^{t},t,\phi(\mathbf{x}_{-L+1:0})))$ (16)
where $\phi(\cdot)$ is the feature extractor. To capture the temporal dependencies across consecutive time steps in time-series data, early approaches such as TimeGrad [35] employ a recurrent neural network (RNN) as the model. This RNN processes historical time-series data and performs predictions in an autoregressive manner. Specifically, TimeGrad leverages the RNN-derived hidden state $\mathbf{h}_{i-1}$ at timestep $i$ to guide denoising. The hidden state $\mathbf{h}_{i-1}$ encapsulates historical dynamics through the feature $\mathbf{h}_{i}=\text{RNN}(\text{concat}(\mathbf{x}_{i},\mathbf{c}_{i}),\mathbf{h}_{i-1})$, where $\mathbf{c}_{i}$ denotes covariates, and $\text{concat}(\cdot)$ represents concatenation operation. Similarly, StochDiff [59] innovates by integrating the diffusion process into the time series modeling stage, applying diffusion at each time step $i$ within a recurrent framework. The model constructs a data-driven prior $\mathbf{z}_{i}\sim\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{h}_{i-1}),\boldsymbol{\Sigma}_{\theta}(\mathbf{h}_{i-1}))$ derived from long short-term memory (LSTM) hidden states $\mathbf{h}_{i-1}$. The hidden state updates as $\mathbf{h}_{i}=\text{LSTM}(\text{concat}(\mathbf{x}_{i},\mathbf{c}_{i}),\mathbf{h}_{i-1})$. This prior refines the reverse diffusion as another conditioning feature. While these recursive approaches capture temporal dynamics to some extent, their autoregressive prediction manner leads to error accumulation, consequently resulting in suboptimal long-term forecasting performance. To better capture long-range temporal dependencies, CSDI [60] replaces RNNs with a transformer-based feature extracting network, utilizing self-attention to establish direct interactions between any two time steps. This network conditions on historical time series $\mathbf{x}_{-L+1:0}$ and outputs predicted time series $\hat{\mathbf{x}}_{1:H}$ for all future steps simultaneously, eliminating error accumulation from autoregressive prediction. Building upon these foundation, SSSD [36] reconstructs the feature extracting network using structured state space models (S4) to enhance efficiency. It employs integrates structured S4 as the backbone of its denoise network architecture, specifically targeting the learning of long-range dependencies within sequences while improving inference speed.
In another line of these approaches, TimeDiff [64] proposes two novel conditioning mechanisms, named future mixup and autoregressive initialization, which differ from prior feature-centric approaches. Future mixup creates a conditioning signal $\mathbf{z}_{\text{mix}}$ during training by stochastically blending a projection of the history $\mathcal{F}(\mathbf{x}_{-L+1:0})$ with elements of the ground-truth future $\mathbf{x}_{1:H}$ according to a mask $\mathbf{m}_{t}$, expressed as
@@
where $\odot$ denotes Hadamard product. Meanwhile, autoregressive initialization produces a simple forecast $\mathbf{z}_{\text{ar}}$ from the history using a pretrained linear model defined by
@@
where $\mathbf{X}^{0}_{i}$ is a matrix formed by replicating the historical value $\mathbf{x}_{i}$ across $H$ future time steps, and $\mathbf{W}_{i}$, $\mathbf{B}$ represent learnable weight and bias parameters. These two components are combined, forming the final condition $\mathbf{c}=\operatorname{concat}\left(\mathbf{z}_{\text{mix}},\mathbf{z}_{\text{ar}}\right)$. These mechanisms strengthen temporal dependence modeling by enriching the conditioning signal and reducing disharmony between the observed past and predicted future segments. Building on TimeDiff’s future mixup conditioning mechanism, mr-Diff [50] extends this concept by incorporating multi-scale historical inputs $\mathbf{X}_{s}$ derived from seasonal-trend decomposition. In addition, Bim-Diff [65] integrates brain-inspired semantic memory $\mathbf{m}_{s}$ and episodic memory $\mathbf{m}_{e}$ as variational priors into the conditioning signal. These memories encode cross-channel recurrent patterns through attention mechanisms:
$\mathbf{m}_{s}^{j}=\sum_{i}\frac{{\rm score}(\mathbf{M}_{s,i},\mathbf{h}_{j})}{\sum_{i}{\rm score}(\mathbf{M}_{s,i},\mathbf{h}_{j})}\mathbf{M}_{s,i},\quad\mathbf{m}_{e}^{j}=\sum_{i}\frac{{\rm score}(\mathbf{M}_{e,i},\mathbf{h}_{j})}{\sum_{i}{\rm score}(\mathbf{M}_{e,i},\mathbf{h}_{j})}\mathbf{M}_{e,i}$ (19)
where $\mathbf{h}_{j}$ represents the query vector for channel $j$. $\mathbf{m}_{s}^{j}$ and $\mathbf{m}_{e}^{j}$ denote the semantic memory and episodic memory on the $j$-th channel, respectively. In addition, $\mathbf{M}_{s}$ and $\mathbf{M}_{e}$ are learnable vectors. The combined memory prior $\mathbf{m}=\mathbf{m}_{s}+\mathbf{m}_{e}$ enriches the condition as $\mathbf{c}_{\text{mix}}=\mathbf{m}_{t}\odot\mathbf{W}\mathbf{m}+(\mathbf{1}-\mathbf{m}_{t})\odot\mathbf{x}_{1:H}$, where $\mathbf{W}$ projects memories into the condition space. This design explicitly incorporates recurring temporal structures into the diffusion process while maintaining mixup’s boundary harmonization. Similarly, to elaborate on the inter-channel modeling approach, CCDM [66] utilizes a channel-aware feature extracting network combining channel-independent multi-layer perceptron encoders (CiDM) for univariate patterns and channel-mixing diffusion Transformers (DiT) for cross-variate correlations. The architecture integrates $d$-channel features through:
$\mathbf{G}=\text{DiT}\left(\text{concat}\left(\{\text{CiDM}(\mathbf{x}_{d,-L+1:0})\}_{d=1}^{D},\mathbf{x}_{1:H}^{t}\right)\right)$ (20)
where $\mathbf{G}$ denotes the fused feature tensor and $D$ is the channel number of time series.
A different line of feature-centric approaches is uncertainty modeling by leveraging specialized feature extraction architectures. For example, D3U [67] leverages a pre-trained point forecasting model to extract deterministic features $\hat{\mathbf{y}}_{1:H}$ from historical data $\mathbf{x}_{-L+1:0}$ and computes residuals $\mathbf{r}_{1:H}^{0}=\mathbf{y}_{1:H}-\hat{\mathbf{y}}_{1:H}$ as high-uncertainty components. These residuals are modeled by a conditional diffusion process where a patch-based denoising network (PatchDN) $\mathbf{r}_{\theta}(\mathbf{r}_{1:H}^{t},t|\mathbf{c})$ predicts $\mathbf{r}_{1:H}^{0}$ at diffusion step $t$ using the reverse transition as
$\mu_{\theta}(\mathbf{r}_{1:H}^{t},t|\mathbf{c})=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{r}_{1:H}^{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\mathbf{r}_{\theta}(\mathbf{r}_{1:H}^{t},t|\mathbf{c})$ (21)
PatchDN utilizes patch embeddings and adaptive layer normalization (AdaLN) to incorporate conditional features $\mathbf{c}$ for scale and shift modulation, thereby refining uncertainty estimation. The final prediction combines deterministic output $\hat{\mathbf{y}}_{1:H}$ and sampled residuals $\hat{\mathbf{r}}_{1:H}^{0}$, yielding $\hat{\mathbf{y}}_{1:H}+\hat{\mathbf{r}}_{1:H}^{0}$. On top of this work, FALDA [54] incorporates Fourier decomposition to decouple time series into distinct non-stationary, stationary, and noise components. The approach utilizes a lightweight denoise network with moving average decomposition and AdaLN to explicitly isolate aleatoric uncertainty patterns. Different from D3U’s transformer-based patch processing, FALDA [54] conditions on pure historical noise terms and targets direct residual reconstruction, separating epistemic uncertainty through component-specific modeling.

### v2

@@ 3 Conditions for TSF Diffusion Models
Based on the conditioning sources for diffusion models in TSF, existing research is categorized into two groups, namely historical time series conditions and multimodal conditions, as illustrated in Fig. 2. Approaches that utilize historical time series conditions generally apply processing techniques from the time series domain to decompose complex temporal dependencies into explicit and simpler features, thereby embedding time series priors into the diffusion model. This feature decomposition reduces the difficulty for the diffusion model to capture temporal correlations and, in certain cases, facilitates more precise uncertainty estimation. The other category applies multimodal conditioning sources, where the core idea is to introduce information from additional modalities to complementarily describe the time series from different perspectives and thus provide richer contextual information. In the following subsections, we elaborate on these two main categories of conditioning sources and their specific implementations in diffusion models.
Figure 3: Conditional sources of the diffusion TSF model. The left panel shows historical and prediction windows, the top-right part illustrates a general diffusion process, and the bottom-right presents condition sources, including historical preprocessing and multimodal inputs. The grey arrows indicate the forward diffusion process, and the red arrows denote the reverse process.
@@ 3.1 Conditioning on Historical Time Series
The historical time series itself directly serves as the conditioning input for the diffusion model without requiring any additional preprocessing, which we denote as Raw [62, 63, 64, 84, 85, 86, 87, 69]. Beyond directly using the raw series, specific time series processing techniques and transformations extract additional information, which injects domain priors as explicit features, as depicted in Fig. 3. According to the employed processing strategy, diffusion-based models conditioning on historical time series are divided into five groups, namely Multi-scale, Decomposition, Spectrum, Retrieval reference, and Latent.
Multi-scale This strategy prepares multiple resolutions of the same time series so that the condition summarizes dynamics from coarse to fine [88, 68]. A generic construction is written as
$\mathbf{x}^{(s)}_{-L+1:0}=\big{(}\mathbf{x}_{-L+1:0}*h^{(s)}\big{)}_{ts},\qquad s\in\{1,\ldots,S\}$ (16)
where $h^{(s)}$ denotes a smoothing kernel at scale $s$, the operator * denotes discrete convolution, the subscript $ts$ denotes subsampling by factor $s$, and $S$ denotes the number of scales. The condition aggregates these resolutions as $\mathbf{c}_{\text{ms}}=\{\mathbf{x}^{(1)}_{-L+1:0},\ldots,\mathbf{x}^{(S)}_{-L+1:0}\}$.
Decomposition This strategy separates the series into interpretable components such as trend, seasonality, and residuals [89, 90]. An additive form is formulated as
$x_{t}=T_{t}+S_{t}+R_{t}$ (17)
where $x_{t}$ denotes the observation at time $t$, $T_{t}$ denotes the trend, $S_{t}$ denotes the seasonal component with period $P$, and $R_{t}$ denotes the remainder. The condition stacks $\mathbf{c}_{\text{dec}}=\{T_{-L+1:0},S_{-L+1:0},R_{-L+1:0}\}$. Spectrum This strategy maps the time series to frequency domain features and combines them with time domain features [66, 91]. A discrete Fourier transform is written as
$\hat{x}_{k}=\sum_{t=0}^{L-1}x_{t}\,e^{-i2\pi kt/L},\qquad k=0,\ldots,L-1$ (18)
where $\hat{x}_{k}$ denotes the complex coefficient at frequency index $k$ and $i$ denotes the imaginary unit. The magnitude and phase are $A_{k}=|\hat{x}_{k}|$ and $\Phi_{k}=\arg(\hat{x}_{k})$. The condition assembles $\mathbf{c}_{\text{spec}}=\{\mathbf{x}_{-L+1:0},A_{0:L-1},\Phi_{0:L-1}\}$ or uses magnitudes only when a phase-free representation is preferred.
Retrieval reference This strategy retrieves similar sequences from a reference set and concatenates them with the target history as auxiliary context [67]. A generic pipeline defines an embedding map $E(\cdot)$ and a cosine similarity $s(\mathbf{u},\mathbf{v})=\frac{\langle E(\mathbf{u}),E(\mathbf{v})\rangle}{|E(\mathbf{u})|_{2},|E(\mathbf{v})|_{2}}$, then selects the top $K$ neighbors $\{\mathbf{u}_{(k)}\}_{k=1}^{K}$ for the query window $\mathbf{x}_{-L+1:0}$. A soft weighting $w_{k}$ is computed by
$w_{k}=\frac{\exp(\tau s_{k})}{\sum_{j=1}^{K}\exp(\tau s_{j})},\qquad s_{k}=s(\mathbf{u}_{(k)},\mathbf{x}_{-L+1:0})$ (19)
with the temperature $\tau>0$. Thus, the reference condition $\mathbf{c}_{\text{ret}}$ is computed through a weighted sum operation by
$\mathbf{c}_{\text{ret}}=\sum_{k=1}^{K}w_{k}\,\mathbf{u}_{(k)}$ (20)
and $\mathbf{c}_{\text{ret}}$ has the same temporal length as each neighbor after alignment or interpolation. Latent This strategy encodes the time series into a compact latent representation that serves as the condition while reducing dimensionality [80]. A variational encoding is formulated by
$q_{\phi}(\mathbf{z}\mid\mathbf{x}_{-L+1:0})=\mathcal{N}\big{(}\boldsymbol{\mu}_{\phi}(\mathbf{x}_{-L+1:0}),\operatorname{diag}(\boldsymbol{\sigma}^{2}_{\phi}(\mathbf{x}_{-L+1:0}))\big{)}$ (21)
with sample $\mathbf{z}\sim q_{\phi}(\mathbf{z}\mid\mathbf{x}_{-L+1:0})$ and condition $\mathbf{c}_{\text{lat}}=\mathbf{z}$, where $\boldsymbol{\mu}_{\phi}(\cdot)$ and $\boldsymbol{\sigma}_{\phi}(\cdot)$ denote encoder outputs parameterized by $\phi$ and $\operatorname{diag}(\cdot)$ denotes a diagonal covariance.
@@ 3.2 Conditioning on Multimodal Data
Multimodal approaches [19, 20, 92, 21] are attracting increasing attention in the TSF domain, and diffusion-based TSF approaches also integrate additional modal information as complementary conditions. Beyond preprocessing and transformation techniques that extract useful information from historical time series as domain-specific priors, multimodal diffusion-based TSF is to incorporate multimodal data as supplementary conditioning sources for the diffusion model, as depicted in Fig. 3. We divide these approaches into Enhanced and External, based on whether the multimodal data originates from the time series itself or from external sources. A unified notation for the multimodal condition $\mathbf{c}_{\mathrm{mm}}$ uses a single formula:
$\mathbf{c}_{\mathrm{mm}}=\{\mathbf{x}_{-L+1:0},\mathbf{u}_{\mathrm{mm}}\}$ (22)
where $\mathbf{u}_{\mathrm{mm}}$ denotes a multimodal embedding produced by a preprocessing map that aggregates auxiliary signals. For Enhanced settings, $\mathbf{u}_{\mathrm{mm}}$ derives from the time series, for example, visual latents from rendered images and a text embedding from a template description. For External settings, $\mathbf{u}_{\mathrm{mm}}$ aggregates aligned side information such as timestamp features and an embedding of external texts.
Enhanced The enhanced multimodal model, for example, LDM4TS [93], derives visual and textual conditions from the same historical time series without using external data. The preprocessing pipeline first converts the series into multi-view images that summarize temporal structure from complementary perspectives, for example, a segmentation-based rearrangement that highlights seasonal stripes, a Gramian angular field that encodes pairwise temporal relations as image intensities, and a recurrence plot that marks repeated state patterns. These images are then passed through a pretrained vision encoder to obtain compact visual latents aligned with the time series. In parallel, the pipeline constructs a template-based textual description that records dataset identity, forecasting task specification, historical and horizon lengths, basic statistics such as min, max, and median, a coarse trend label, and top autocorrelation lags. Then, a frozen language model encodes this description into a text embedding. Therefore, the final conditioning set contains visual latents and a text vector, both deterministically computed from the input time series.
External The external multimodal approach, for example, MCD-TSF [94] utilizes side information aligned with the time axis, including timestamp metadata and news text with the target series. Preprocessing firstly normalizes calendar attributes, including day of week, month, and holiday flags, into fixed-length timestamp features and aligns them to each historical step. It then collects documents that fall within the same interval as the history window, performs standard text cleaning and deduplication, and encodes each document with a pretrained language model. The resulting vectors are aggregated to form a single text representation. The multimodal condition is obtained by aligning the numeric history with the timestamp features and the aggregated text vector. When text is missing, the conditioning reduces to numeric history plus timestamps, which keeps the pipeline applicable in partially observed settings.
@@ 4 Condition Integration
Figure 4: Condition integration approaches for diffusion-based TSF models. The left part shows the historical and prediction windows. The bottom right part denotes a generic condition source. The top right part illustrates that the condition firstly enters a denoise network (e.g., a recurrent, convolutional, attention-based, or MLP-based structure) to extract informative features. These features are then incorporated into the diffusion process, which is organized into two categories referred to as enhanced process and reformulated process, representing alternative ways of embedding conditioning information into the generative trajectory.
In addition to the source used for conditioning, diffusion-based TSF approaches also differ in their approaches to integrating conditioning information into the generative diffusion process. Existing studies integrate time series conditions by designing a specialized denoising network for the diffusion model or modifying the diffusion process. Therefore, we classify existing condition integration mechanisms into two primary categories, feature-centric and diffusion-centric approaches, as depicted in Fig. 4. Feature-centric strategies [62, 86, 63] generally employ the standard diffusion model but rely on a well-designed denoise network to extract diffusion-step-aware features from the historical conditioning data. These features then act as input to the learned denoising function, guiding the reverse process. Diffusion-centric approaches [95, 96, 97] propose modifications to the diffusion process. Such modifications incorporate historical data priors directly into both forward and reverse diffusion trajectories, enabling finer-grained control over the generation. The following subsections explain the various studies, detailing their methodologies and characteristics.
@@ 4.1 Feature-centric Condition
Feature-centric diffusion approaches retain the standard forward noising and reverse denoising processes while placing the design focus on the denoising network as a feature extractor that maps the historical window into conditioning features for every diffusion step. A generic reverse transition for the future horizon $\mathbf{x}_{1:H}$ at step $t$ writes
$\mathbf{x}_{1:H}^{t-1}\sim\mathcal{N}\big{(}\boldsymbol{\mu}_{\theta}(\mathbf{x}_{1:H}^{t},t,\phi(\mathbf{x}_{-L+1:0})),\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{1:H}^{t},t,\phi(\mathbf{x}_{-L+1:0}))\big{)}$ (23)
The function $\phi(\cdot)$ denotes the feature extractor. Based on the core model structure of the feature extractor, we categorize feature-centric approaches as MLP-based, Recurrent, Convolutional, and Attention-based.
MLP-based feature extractors MLP-based approaches [89, 69, 91] use feed-forward blocks to distill patterns from the history and to modulate conditioning through adaptive layer normalization, which is formulated as
$\operatorname{AdaLN}(\mathbf{h},\mathbf{c})=\gamma(\mathbf{c})\odot\frac{\mathbf{h}-\mu(\mathbf{h})}{\sigma(\mathbf{h})}+\beta(\mathbf{c})$ (24)
where $\operatorname{AdaLN}(\mathbf{h},\mathbf{c})$ denotes adaptive layer normalization of $\mathbf{h}$ conditioned on $\mathbf{c}$, $\mathbf{h}$ denotes an intermediate feature, $\mathbf{c}$ denotes conditioning features from the history, $\mu(\mathbf{h})$ and $\sigma(\mathbf{h})$ denote feature wise mean and standard deviation, $\odot$ denotes Hadamard product, and $\gamma(\cdot)$ and $\beta(\cdot)$ are MLP outputs driven by $\mathbf{c}$. FALDA [91] performs a light decomposition that splits the signal as $x_{t}=T_{t}+S_{t}+N_{t}$ with a moving average trend $T_{t}$, a frequency-selected stationary part $S_{t}$, and the residual noise $N_{t}=x_{t}-T_{t}-S_{t}$. The decomposed features drive an MLP path under adaptive layer normalization with scale and shift produced from the condition. D3U [69] prepares a deterministic forecast $\hat{\mathbf{y}}_{1:H}$ and models residual uncertainty $\mathbf{r}_{1:H}^{0}=\mathbf{y}_{1:H}-\hat{\mathbf{y}}_{1\>H}$ with an MLP based conditional predictor under AdaLN.
Recurrent feature extractors Recurrent approaches [62, 86] leverage hidden state dynamics to propagate historical context through time and across diffusion steps. TimeGrad [62] processes the observed sequence with a recurrent cell and maintains a hidden state that updates as
$\mathbf{h}_{i}=\operatorname{RNN}(\operatorname{concat}(\mathbf{x}_{i},\mathbf{c}_{i}),\mathbf{h}_{i-1})$ (25)
where $\mathbf{x}_{i}$ denotes the value at time index $i$, $\mathbf{c}_{i}$ denotes covariates at index $i$, and $\operatorname{concat}(\cdot,\cdot)$ concatenates its arguments along the feature (channel) dimension to form the RNN input at step $i$. The denoiser at horizon index $i$ then conditions on $\mathbf{h}_{i-1}$. StochDiff [86] integrates an LSTM into the feature extractor and defines a data-driven prior at each horizon index, which is formulated as
$\mathbf{z}_{i}\sim\mathcal{N}\big{(}\boldsymbol{\mu}_{\theta}(\mathbf{h}_{i-1}),\boldsymbol{\Sigma}_{\theta}(\mathbf{h}_{i-1})\big{)}$ (26)
$\mathbf{h}_{i}=\operatorname{LSTM}(\operatorname{concat}(\mathbf{x}_{i},\mathbf{c}_{i}),\mathbf{h}_{i-1})$ (27)
where $\mathbf{z}_{i}$ serves as a prior feature for the reverse step of index $i$ and depends on the previous hidden state $\mathbf{h}_{i-1}$. This design supplies inductive bias for sequential order and regime shifts while remaining compatible with non-autoregressive denoising over the full horizon.
Convolutional feature extractors Convolutional approaches [64, 84, 93] rely on one-dimensional convolutions and convolutional sequence kernels to capture temporal dependencies with high efficiency. SSSD [64] adopts an S4 based convolutional backbone where a learned kernel $K$ induces a causal convolution
$\mathbf{y}=(K*\mathbf{z}^{t})=\sum_{\tau\geq 0}K[\tau]\mathbf{z}^{t-\tau}$ (28)
where $\mathbf{z}^{t}$ denotes the diffused input at step $t$, $K[\tau]$ denotes the discrete kernel produced by the state space parameterization, and the sum runs over nonnegative lags. TimeDiff [84] employs an encoder and a decoder made of stacked one-dimensional convolutions together with training time conditioning that aligns the boundary between past and future. The future mixup condition uses a stochastic mask $\mathbf{m}_{t}$ and a projection $\mathcal{F}$ to form
@@
where $\mathbf{m}_{t}$ samples a binary mask at diffusion step $t$ and $\mathcal{F}$ denotes a learnable projection of the history into the horizon shape. The linear initialization $\mathbf{z}_{\text{ar}}$ is computed by
@@
Here $\mathbf{X}^{0}_{i}$ replicates $\mathbf{x}_{i}$ across $H$ future positions and $\mathbf{W}_{i}$ and $\mathbf{B}$ denote learnable weights and bias. The final condition uses $\mathbf{c}=\operatorname{concat}(\mathbf{z}_{\text{mix}},\mathbf{z}_{\text{ar}})$. LDM4TS [93] converts the series into multi-view images and applies latent diffusion with a convolutional UNet denoiser on image latents. A compact notation writes $\mathbf{z}_{\text{img}}=E(\mathcal{I}(\mathbf{x}_{-L+1:0}))$ where $\mathcal{I}(\cdot)$ denotes the image transformation and $E(\cdot)$ denotes the vision encoder.
Attention-based feature extractors Attention-based approaches [63, 66, 80, 85, 67, 87, 94] use self-attention to model long-range dependencies over time and across variables and to integrate side information or memories within the denoiser. CSDI [63] stacks temporal and feature-wise attention to produce conditioning features for imputation and forecasting with
$\operatorname{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{softmax}\!\big{(}\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d_{k}}\big{)}\mathbf{V}$ (31)
$\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q},\ \mathbf{K}=\mathbf{X}\mathbf{W}_{K},\ \mathbf{V}=\mathbf{X}\mathbf{W}_{V}$ (32)
where $\mathbf{Q},\mathbf{K},\mathbf{V}$ are the query, key, and value matrices obtained from the token matrix $\mathbf{X}$ via learned projections $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$. $\operatorname{Attn}(\cdot)$ denotes scaled dot-product attention, and $\operatorname{softmax}(\cdot)$ performs row-wise normalization to convert similarities into weights over keys for each query. $\mathbf{X}$ stacks temporal and variable tokens, and $d_{k}$ is the key dimension. Diffusion-TS [66] adopts a Transformer encoder and decoder with decomposition guided representations and augments the diffusion objective with a frequency loss $\mathcal{L}_{\text{freq}}$:
$\mathcal{L}_{\text{freq}}=\big{\|}\mathcal{F}(\hat{\mathbf{x}}_{1:H})-\mathcal{F}(\mathbf{x}_{1:H})\big{\|}^{2}$ (33)
where $\mathcal{F}(\cdot)$ denotes the discrete Fourier transform and $\hat{\mathbf{x}}_{1:H}$ denotes the model output. LDT [80] performs diffusion in a compact latent space with a Transformer denoiser and self-conditioning, where the latent at the step $t$ augments the input for the step t1t-1 through
$\tilde{\mathbf{z}}^{t-1}=\operatorname{concat}\big{(}\mathbf{z}^{t-1},g(\mathbf{z}^{t})\big{)}$ (34)
where $\tilde{\mathbf{z}}^{t-1}$ denotes the self-conditioned latent input to the denoiser at step t1t-1 formed by concatenating $\mathbf{z}^{t-1}$ with the projected latent $g(\mathbf{z}^{t})$, $\mathbf{z}^{t}$ denotes the latent variables at step $t$, and $g(\cdot)$ denotes a learned projection for self-conditioning. Bim-Diff [85] introduces semantic and episodic memories $\mathbf{M}_{s}$ and $\mathbf{M}_{e}$ that produce memory priors through attention
$\mathbf{m}_{s}^{j}=\sum_{i}\omega_{s,i}^{j}\mathbf{M}_{s,i},\qquad\mathbf{m}_{e}^{j}=\sum_{i}\omega_{e,i}^{j}\mathbf{M}_{e,i}$ (35)
where the index $j$ denotes the channel index, $\omega_{s,i}^{j}$ and $\omega_{e,i}^{j}$ denote softmax normalized scores computed from queries $\mathbf{h}_{j}$, and $\mathbf{M}_{s,i}$ and $\mathbf{M}_{e,i}$ denote memory items. The combined prior $\mathbf{m}=\mathbf{m}_{s}+\mathbf{m}_{e}$ passes through a projection and a mask to yield mixed condition $\mathbf{c}_{\text{mix}}$:
$\mathbf{c}_{\text{mix}}=\mathbf{m}_{t}\odot\mathbf{W}\mathbf{m}+(\mathbf{1}-\mathbf{m}_{t})\odot\mathbf{x}_{1:H}$ (36)
where $\mathbf{m}_{t}$ denotes a stochastic mask and $\mathbf{W}$ denotes a learnable projection that maps the memory prior into the condition space. RATD [67] retrieves neighbor time series $\mathbf{u}_{k}$} from a reference set and injects them by cross attention $\operatorname{Attn(\mathbf{Q},\mathbf{K},\mathbf{V})}$ with $\mathbf{Q}$ from the current hidden representation and with $\mathbf{K}$ and $\mathbf{V}$ from the retrieved set. CCDM [87] combines per-channel MLP encoders with a channel-mixing diffusion Transformer, where a channel-independent dense encoder produces univariate features and a mixing block computes the channel-mixed representation $\mathbf{G}$ by
$\mathbf{G}=\operatorname{DiT}\big{(}\operatorname{concat}(\{\mathbf{f}_{d}\}_{d=1}^{D},\mathbf{x}_{1:H}^{t})\big{)}$ (37)
where $\mathbf{G}$ is subsequently fed to the output decoder for step-wise noise regression, $D$ is the number of channels, $\mathbf{f}_{d}=\operatorname{MLP}(\mathbf{x}_{d,-L+1:0})$ is the per-channel embedding, and $\operatorname{DiT}(\cdot)$ denotes a diffusion Transformer block that aggregates cross-channel temporal modes via multi-head self-attention. MCD-TSF [94] incorporates multimodal conditions such as text or spatial descriptors and fuses them with the historical series through cross-attention modules that generate conditioning features for the denoiser. A compact formulation reuses the cross-attention operator above with queries from the time series tokens and keys and values from multimodal tokens.